OpenStudy (anonymous):

To write 1014 in standard form, you would use: a 10 and fourteen 0s. a 1 and fourteen 0s. fourteen 10s. a 14 and ten 10s.

OpenStudy (cwrw238):

i dont understand the question do you mean 10^14 ?

OpenStudy (anonymous):

Me Eithier

OpenStudy (anonymous):

idk

OpenStudy (cwrw238):

i think it probably is - thats the only thing which goes with the answers if so the answer is 1 and fourteen zeros

OpenStudy (cwrw238):

since 10^1 = 10 10^2 = 100 10^3 = 1000 etc